OpenStudy (anonymous):

Can someone please tell me if I am doing this correct so far please. Integration: I will type it below

OpenStudy (anonymous):

sure go ahead

OpenStudy (anonymous):

\[\int\limits (x^2+1 )e^x dx\] let u=x^2 +1 dv = e^x dx du = 2x dx v=e^x dx then following the formula I put \[\int\limits (x^2+1)e^x = x^2(e^x) - \int\limits (e^x)(2x dx)\]

OpenStudy (anonymous):

make your live easy and multiply out first

OpenStudy (anonymous):

\[\int x^2e^xdx+\int e^xdx\] second integral is just \[e^x\] and parts for the first one

OpenStudy (kinggeorge):

Also, if you do decide to use \(u=x^2+1\), then the term in front should be \[(x^2+1)(e^x)\]not \[x^2(e^x)\]

OpenStudy (anonymous):

always when solving integrals, it becomes easy when you expand the factorized stuff

OpenStudy (anonymous):

actuall you will need parts twice

OpenStudy (anonymous):

yeah I just wanted to make sure I had the first part correct. for my second part I made u=x du=dx dv = e^x and v=e^x

OpenStudy (anonymous):

how do I expand it Shana?

OpenStudy (anonymous):

take 2 out of the integral your u n v are correct. go ahead with that. wait i ll attach a drawing..

OpenStudy (anonymous):